OpenStudy (anonymous):

a = (2,-1,3), b = (7,5,1), c = (5,5,0) Determin an equation of the plane containing the points a, b, and c

OpenStudy (anonymous):

well the equation: -2x + 3y + 4z =5 i did it by finding a normal vector to ab , ac you can do this by the determinant : ab : (5,6,-2) ac (3,6,-3) determinant : i j k 5 6 -2 -> (ab) 3 6 -3 -> (ac) the normal vector from this determinant : (-6,9,12) then i did dot product between (x-2,y+1,z-3) which is any vector through the point a and the normal (-6,9,12) and the dot product must be equal to zero since its dot product with a normal vector. (x-2,y+1,z-3)(-6,9,12) -> -2x + 3y + 4z =5

OpenStudy (anonymous):

\[-6x+9y+12z=15\] If you would like me to explain just ask.

OpenStudy (anonymous):

hmm

OpenStudy (amistre64):

the negative in front is seen as a fauxpaux; and the unreduced form is bad posture as well

OpenStudy (amistre64):

zero out a point to see 2 vectors on the plane (2,-1,3) (7,5,1) (5,5,0) -5-50 -5-50 -5-50 -------------------- -3,-6, 3 2,0,1 <1,2,-1> and <2,0,1> are on the plane

OpenStudy (amistre64):

cross them to get the normal: <1,2,-1> <2, 0, 1> --------- Nx = 2- 0 = 2 Ny = -2-1 = -3 Nz = 0 - 4 = -4 N = <2,-3,-4> right?

OpenStudy (amistre64):

Now pick any given point (Px,Py,Pz) and stick it into the equation 2(x-Px)-3(y-Py)-4(z-Pz) = 0

OpenStudy (amistre64):

(5,5,0) 2(x-5)-3(y-5)-4(z-0) = 0 2x-10-3y+15-4z = 0 2x -3y -4z +5 = 0

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I need to digest that a bit, but I should be able to get it

OpenStudy (anonymous):

Thanks!